OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 2y = 6 x + 3y = −1 a. −2 b. 2 c. 5 d. −5

OpenStudy (anonymous):

Can someone please walk me through how to do this

OpenStudy (solomonzelman):

first equation times 3 and second equation times -2 (2x + 2y = 6) × 3 (x + 3y = −1) × (-2) this will get you a new (but equivalent to the previous) system of equations. 6x + 6y = 18 -2x - 6y = 2 now add the equations and solve for x

OpenStudy (anonymous):

So you just take one of the equations and multiply it to the other equation?

OpenStudy (triciaal):

first equation divide each term by 2

OpenStudy (anonymous):

Would it be x+y=6 then?

OpenStudy (solomonzelman):

I didnàt multiply equation times equation, all I did is that I multiplied each term in the equation times some number.... i did this for both of the equations, and choose the numbers by which to multiply in a way that would eliminate the y (since the variable you need is x).

OpenStudy (anonymous):

So you would come up with 6x+6y= 18? for the first equation and -2x+-6=2?

OpenStudy (solomonzelman):

-2x - 6y = 2 for the second equation, you meant (right ?)

OpenStudy (anonymous):

Sorry yes thats what I meant

OpenStudy (solomonzelman):

yes, that is what you inded come up with.

OpenStudy (solomonzelman):

itàs ok...)

OpenStudy (solomonzelman):

Now, add the equations

OpenStudy (solomonzelman):

6x + 6y = 18 -2x - 6y = 2

OpenStudy (anonymous):

Okay so it would be 4x + 18y = 20?

OpenStudy (anonymous):

Oh wait 12y*

OpenStudy (solomonzelman):

18y?

OpenStudy (solomonzelman):

no, not 12. when you add 6 and -6 you get 6+(-6)=6-6=0 and so when you add 6y and -6y you get (6y)+(-6y)=??

OpenStudy (anonymous):

Oh so it would be 4x = 20?

OpenStudy (solomonzelman):

so x is?

OpenStudy (anonymous):

5?

OpenStudy (solomonzelman):

yes, that is correct!

OpenStudy (anonymous):

Thank you so much! That makes a lot more sense now!

OpenStudy (solomonzelman):

very nice to hear that... you are welcome!

OpenStudy (triciaal):